OpenStudy (anonymous):

Solve the absolute value inequality: |4x + 3| >= 19 i know the answer but i dont no the difference between or and and

OpenStudy (anonymous):

4x+3>=19 -4x-3>=19 4x>=16 -4x>=22 x>=4 x<=-22/4

OpenStudy (anonymous):

x <= -5.5 or x>= 4 is the answer but wat will go in middle and or or

OpenStudy (anonymous):

there are two same answers and in one it has or and in the other it has and

OpenStudy (apoorvk):

okay, we proceed in the same way as earlier! Since *absolute value* of "4x+3" is greater than or equal to 19, we conclude that 4x+3 is greater than 19, or lesser than 'negative-19'. So, \(4x+3\ge 19\) AND \(4x+3\le-19\) Can you solve these two inequalities?

OpenStudy (anonymous):

yeah u get the answer above

OpenStudy (anonymous):

i know how to solve i just dont no when to use and and when to use or

OpenStudy (anonymous):

in the middle of both equations do i put and or a or